solve by completing the square
x^2-6x-4=0
simplify your answer. type an exact answer using radicals as needed. use integers or fractions for any numbers in the expression. use a comma to seperate answers as needed.

Answer by mananth(16946)

x^2-6x-4=0
x^2-6x-4+13=0+13
x^2+6x+9=13
(x+3)^2=13
Take the squareroot
x+3= -+sqrt(13)
x=3+sqrt(13)
OR
x=3-sqrt(13)